OpenStudy (nthenic_oftime):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work

OpenStudy (calculusxy):

Do you know the quadratic formula?

OpenStudy (calculusxy):

Quadratic Formula: \(\huge x =\frac{-b \pm \sqrt{(b)^2 - 4ac}}{2a}\) Standard Form of Quadratics: \(\large ax^2 + bx + c = 0\) (where \(a\) and \(b\) are coefficients and \(c\) is a constant).

OpenStudy (calculusxy):

It would be best to identify the \(a\), \(b\), and \(c\) terms in your quadratic equation \(6x^2 + 12x + 5 = 0\).

OpenStudy (calculusxy):

@nthenic_oftime

OpenStudy (mathmale):

Agreed. The coefficients are a=6, b=? and c=?